(You are hoping to buy a house in the future and recently received an inheritance of ​$22,000.You intend to use your inheritance as a down payment on your house.

a.  If you put your inheritance in an account that earns 9 percent interest compounded​ annually, how many years will it be before your inheritance grows to ​$35,000​?

b.  If you let your money grow for 9.5 years at 9 percent​, how much will you​ have?

c.  How long will it take your money to grow to $35,000 if you move it into an account that pays 3 percent compounded​ annually? How long will it take your money to grow to​$35,000 if you move it into an account that pays 12 percent​?

d.  What does all this tell you about the relationship among interest​ rates, time, and future​ sums?
Answer #1

a.  If you put your inheritance in an account that earns 9 percent interest compounded​ annually, how many years will it be before your inheritance grows to ​$35,000​?
Using financial calculator
PV=-22000
PMT=0
FV=35000
I/Y=9%
CPT N=5.387770019 years

b.  If you let your money grow for 9.5 years at 9 percent​, how much will you​ have?
Using financial calculator
N=9.5
I/Y=9%
PV=-22000
PMT=0
CPT FV=49885.51

c.  How long will it take your money to grow to $35,000 if you move it into an account that pays 3 percent compounded​ annually?
Using financial calculator
I/Y=3%
PMT=0
FV=35000
PV=-22000
CPT N=15.70786273 years

How long will it take your money to grow to​$35,000 if you move it into an account that pays 12 percent​?
Using financial calculator
I/Y=12%
FV=35000
PV=-22000
PMT=0
CPT N=4.096982215 years

d.  What does all this tell you about the relationship among interest​ rates, time, and future​ sums?
Higher the rate higher is the future sum
Higher the time higher is the future sum